# STAR: A Benchmark for Astronomical Star Fields Super-Resolution

### v1

@@ 2 Related Works
shan2025galaxy 9383 Galaxy $\times$ $\times$ $\times$
song2024improving 1597 Solar $\checkmark$ $\times$ $\times$
@@
luo2025cross 17000 Galaxy $\times$ $\times$ $\times$
li2022self 14604 Galaxy $\checkmark$ $\times$ $\times$
STAR 54738 Star field $\checkmark$ $\checkmark$ $\checkmark$
@@ 4.4 Flux Consistency Loss
The idea of the Flux Consistency Loss is to train the model to generate flux consistent prediction. We propose a simple yet effective scheme to achieve this goal by highlighting flux-related regions. We first operate the flux map generation scheme described in Section 4.2 on the ground-truth image, denoted as $M$. They we use this map to weighted the pixel wised supervision as follows:
@@ 5.3 Ablation Study
To evaluate the contribution of each proposed component in our FISR framework, we conduct comprehensive ablation studies on the STAR dataset under the ×$\times 2$ super-resolution setting. Tab. 3 presents the performance changes when incorporating or removing the Flux Guidance Generation (FGG), Flux Guidance Controller (FGC) and Flux Consistency Loss (FCL).
@@
Table 2: Performance comparison of different methods under ×$\times 2$ and ×$\times 4$ super-resolution. Evaluation metrics include PSNR, SSIM, and Flux Error.
@@
Table 4: Comparison of different methods with and without FCL under ×$\times 2$ and ×$\times 4$ ASR.
@@ 7 Limitations and future work
While our study offers promising insights, it has a few limitations that merit further exploration. First, our experiments are based on observations from a single telescope, the HST WFC/ACS with the F814W filter, which may limit the generalizability of our findings to other instruments or observational contexts. Additionally, although our network design performs well, it could benefit from incorporating more domain-specific optimizations rooted in astronomical knowledge, such as leveraging physical principles or astronomical priors to enhance performance in complex scenarios. These areas present opportunities for future refinement. Looking forward, we aim to broaden the applicability of our method by extending it to a wider array of advanced telescopes, such as the James Webb Space Telescope (JWST) pontoppidan2022jwst or the upcoming Large Synoptic Survey Telescope (LSST) ivezic2019lsst , to explore its potential across diverse astronomical contexts. Furthermore, we plan to enhance our network design by integrating more astronomy-driven optimizations, incorporating physical knowledge and astronomical priors to better address challenges like crowded stellar regions or variable noise conditions. Through these efforts, we hope to make modest contributions to the field of astronomical image processing, fostering the development of more robust and adaptable tools for future discoveries.
@@ 8 Acknowledgments
This work is supported by the Shanghai Artificial Intelligence Laboratory. We sincerely thank Hao Du, Yating Liu, Jiaze Li, Yinfan Hua, Jun Yao for their invaluable guidance and insightful feedback throughout this research.
@@ Appendix B PSF Details
We simulate the imaging blur in the STAR dataset using two PSF models: the Gaussian PSF and the Airy PSF [gardill2022super], aiming to increase training data diversity. The Gaussian PSF is a simple model often used to approximate blur in astronomical observations [74, 75]. In contrast, the Airy PSF captures diffraction effects from a telescope’s circular aperture, making it suitable for space-based instruments like HST [24].
@@ Appendix C Additional Experiments with Gaussian + Airy PSFs
We compare the performance of different methods under ×$\times 2$ super-resolution with Gaussian PSF and Airy PSF on the STAR dataset, analyzing the results model-wise and loss-wise. Tab. 5 compares the performance of all methods in this setting. FIRS surpasses baselines like SwinIR and RCAN, achieving a 3.05% higher PSNR and 26.45% lower FE than SwinIR, demonstrating its superior ability to recover fine stellar details and preserve flux accuracy in astronomical image super-resolution. Additionally, Tab. 6 compares EDSR, RCAN, and SwinIR with and without FCL to focus on the impact of FCL across baseline methods. For instance, SwinIR with FCL improves PSNR by 1.27% and reduces FE by 10.88% compared to the version without FCL, while RCAN with FCL improves PSNR by 1.14% and reduces FE by 20.63%, highlighting FCL’s role in enhancing image quality and flux preservation.
@@ Appendix D Additional visualizations
We present additional visualizations to demonstrate the effectiveness of our approach in star-field super-resolution (ASR) tasks. Fig. 6 displays the ×$\times 2$ super-resolution results for the Gaussian PSF experiment, comparing baselines (EDSR, RCAN, PromptIR, SwinIR, HAT) against our FIRS model. The visualizations reveal that FIRS consistently outperforms all baselines, achieving superior visual quality with finer stellar details and sharper structures. To further quantify these improvements, we compute the KL divergence and JS divergence between the intensity distributions of the predicted and ground truth values in selected regions, following the experimental settings in the original submission. The results show that FIRS significantly reduces distribution discrepancies compared to SwinIR and HAT, confirming its superior capability in preserving stellar details and flux accuracy in ASR tasks.
@@ Appendix E Hyperparameters Tuning
We tune the parameter $\lambda$ to balance the Flux-Consistent Loss (FCL) and reconstruction loss in our star-field super-resolution (ASR) model. We evaluate different $\lambda$ values (0.1, 0.05, and 0.01) under the ×$\times 2$ Gaussian PSF + Airy PSF setting, with results shown in Tab. 9. The performance metrics show that $\lambda=0.01$ yields the best results, improving PSNR by 1.40% and reducing FE by 15.88% compared to $\lambda=0.1$. These results indicate that a proper $\lambda$ matters in the balance between reconstruction loss and the Flux-Consistent Loss. Fortunately, 0.01 seems to perform well in most cases.
@@ Appendix F Experimental setting/details
Appendix G Limitations and future work
@@
Table 9: Ablation study on the penalty factor $\lambda$ (×$\times 2$ on Gaussian PSF + Airy PSF).

### v2

@@ 2 Related Works
QQ Shan et al shan2025galaxy 9383 Galaxy $\times$ $\times$ $\times$
W Song et al song2024improving 1597 Solar $\checkmark$ $\times$ $\times$
@@
ZJ Luo et al luo2025cross 17000 Galaxy $\times$ $\times$ $\times$
WJ Li et al li2022self 14604 Galaxy $\checkmark$ $\times$ $\times$
[…]
@@ 4.4 Flux Consistency Loss
The idea of the Flux Consistency Loss is to train the model to generate flux consistent prediction. We propose a simple yet effective scheme to achieve this goal by highlighting flux-related regions. We first operate the flux map generation scheme described in Section 4.2 on the ground-truth image, denoted as $M$. Then we use this map to weighted the pixel wised supervision as follows:
@@ 5.3 Ablation Study
To evaluate the contribution of each proposed component in our FISR framework, we conduct comprehensive ablation studies on the STAR dataset under the $\times 2$ super-resolution setting. Tab. 3 presents the performance changes when incorporating or removing the Flux Guidance Generation (FGG), Flux Guidance Controller (FGC) and Flux Consistency Loss (FCL).
@@
Table 2: Performance comparison of different methods under $\times 2$ and $\times 4$ super-resolution. Evaluation metrics include PSNR, SSIM, and Flux Error.
@@
Table 4: Comparison of different methods with and without FCL under $\times 2$ and $\times 4$ ASR.
@@ 5.4 Quality Analysis
5.5 Evaluation on Downstream Scientific Tasks
To address concerns about error propagation to scientific outputs, we evaluated our method on two representative downstream tasks: stellar mass estimation and weak lensing shear measurement.
For stellar mass estimation, we applied a simplified photometric pipeline to the STAR test set, converting predicted fluxes to magnitudes ($\text{mag}=-2.5\times\log_{10}(\text{flux})+25.0$) and inferring stellar masses using a constant mass-to-light ratio ($M/L\approx 3.0$). As shown in Table 5, our FISR method achieves the lowest predicted magnitude error ($1.66\times 10^{-7}$) alongside RealESRGAN ($1.67\times 10^{-7}$), and demonstrates the best stellar mass MAE ($2.81\times 10^{-8}$), significantly outperforming other methods including EDSR ($1.37\times 10^{-7}$) and Bicubic ($1.79\times 10^{-7}$).
For weak lensing shear measurement, a critical task for cosmological studies, we computed shear components $\gamma=\gamma_{1}+i\gamma_{2}$ from SEP-detected galaxy ellipses, where:
$\gamma_{1}$ $=\frac{a^{2}-b^{2}}{a^{2}+b^{2}}\times\cos(2\theta)$ (6)
$\gamma_{2}$ $=\frac{a^{2}-b^{2}}{a^{2}+b^{2}}\times\sin(2\theta)$ (7)
with semi-major axis $a$, semi-minor axis $b$, and position angle $\theta$. We measured the mean shear difference $|\gamma_{\text{pred}}-\gamma_{\text{gt}}|$ between predictions and ground truth. The results in Table 5 show that FISR achieves competitive shear preservation ($1.87\times 10^{-1}$), comparable to the best-performing HAT ($1.88\times 10^{-1}$) and superior to methods like RCAN ($2.14\times 10^{-1}$). These comprehensive evaluations demonstrate that our flux-preserving approach maintains both photometric accuracy and morphological fidelity essential for downstream scientific inference, validating its practical utility in astronomical applications.
Table 5: Evaluation on downstream scientific tasks. We report predicted magnitude error, stellar mass MAE, and mean shear error for different methods. Lower values indicate better performance.
Method Pred. Mag. Error Mass MAE Mean Shear Error
Bicubic $3.23\times 10^{-7}$ $1.79\times 10^{-7}$ $2.10\times 10^{-1}$
SwinIR $2.02\times 10^{-7}$ $5.19\times 10^{-8}$ $1.98\times 10^{-1}$
EDSR $2.96\times 10^{-7}$ $1.37\times 10^{-7}$ $2.14\times 10^{-1}$
RCAN $2.01\times 10^{-7}$ $5.36\times 10^{-8}$ $2.06\times 10^{-1}$
HAT $1.67\times 10^{-7}$ $3.06\times 10^{-8}$ $1.88\times 10^{-1}$
RealESRGAN $3.37\times 10^{-7}$ $3.99\times 10^{-7}$ $1.95\times 10^{-1}$
FISR (Ours) $\mathbf{1.66\times 10^{-7}}$ $\mathbf{2.81\times 10^{-8}}$ $\mathbf{1.87\times 10^{-1}}$
@@ 7 Limitations and future work
While our study offers promising insights, it has a few limitations that merit further exploration. First, our experiments are based on observations from a single telescope, the HST WFC/ACS with the F814W filter, which may limit the generalizability of our findings to other instruments or observational contexts. Additionally, although our network design performs well, it could benefit from incorporating more domain-specific optimizations rooted in astronomical knowledge, such as leveraging physical principles or astronomical priors to enhance performance in complex scenarios. These areas present opportunities for future refinement. Looking forward, we aim to broaden the applicability of our method by extending it to a wider array of advanced telescopes, such as the James Webb Space Telescope (JWST) pontoppidan2022jwst or the upcoming Large Synoptic Survey Telescope (LSST) ivezic2019lsst , to explore its potential across diverse astronomical contexts. Through these efforts, we hope to make modest contributions to the field of astronomical image processing, fostering the development of more robust and adaptable tools for future discoveries.
@@ 8 Acknowledgments
This work was supported by the Shanghai Artificial Intelligence Laboratory. This work was also supported by the JC STEM Lab of AI for Science and Engineering, funded by The Hong Kong Jockey Club Charities Trust, and the Research Grants Council of Hong Kong (Project No. CUHK14213224). We sincerely thank Hao Du, Yating Liu, Jiaze Li, Yingfan Hua, and Jun Yao for their invaluable guidance and insightful feedback throughout this research.
@@ Appendix B PSF Details
We simulate the imaging blur in the STAR dataset using two PSF models: the Gaussian PSF and the Airy PSF [74], aiming to increase training data diversity. The Gaussian PSF is a simple model often used to approximate blur in astronomical observations [75, 76]. In contrast, the Airy PSF captures diffraction effects from a telescope’s circular aperture, making it suitable for space-based instruments like HST [24].
@@ Appendix C Additional Experiments with Gaussian + Airy PSFs
We compare the performance of different methods under $\times 2$ super-resolution with Gaussian PSF and Airy PSF on the STAR dataset, analyzing the results model-wise and loss-wise. Tab. 6 compares the performance of all methods in this setting. FIRS surpasses baselines like SwinIR and RCAN, achieving a 3.05% higher PSNR and 26.45% lower FE than SwinIR, demonstrating its superior ability to recover fine stellar details and preserve flux accuracy in astronomical image super-resolution. Additionally, Tab. 7 compares EDSR, RCAN, and SwinIR with and without FCL to focus on the impact of FCL across baseline methods. For instance, SwinIR with FCL improves PSNR by 1.27% and reduces FE by 10.88% compared to the version without FCL, while RCAN with FCL improves PSNR by 1.14% and reduces FE by 20.63%, highlighting FCL’s role in enhancing image quality and flux preservation.
@@ Appendix D Additional visualizations
We present additional visualizations to demonstrate the effectiveness of our approach in star-field super-resolution (ASR) tasks. Fig. 6 displays the $\times 2$ super-resolution results for the Gaussian PSF experiment, comparing baselines (EDSR, RCAN, PromptIR, SwinIR, HAT) against our FIRS model. The visualizations reveal that FIRS consistently outperforms all baselines, achieving superior visual quality with finer stellar details and sharper structures. To further quantify these improvements, we compute the KL divergence and JS divergence between the intensity distributions of the predicted and ground truth values in selected regions, following the experimental settings in the original submission. The results show that FIRS significantly reduces distribution discrepancies compared to SwinIR and HAT, confirming its superior capability in preserving stellar details and flux accuracy in ASR tasks.
@@ Appendix E Hyperparameters Tuning
We tune the parameter $\lambda$ to balance the Flux-Consistent Loss (FCL) and reconstruction loss in our star-field super-resolution (ASR) model. We evaluate different $\lambda$ values (0.1, 0.05, and 0.01) under the $\times 2$ Gaussian PSF + Airy PSF setting, with results shown in Tab. 17. The performance metrics show that $\lambda=0.01$ yields the best results, improving PSNR by 1.40% and reducing FE by 15.88% compared to $\lambda=0.1$. These results indicate that a proper $\lambda$ matters in the balance between reconstruction loss and the Flux-Consistent Loss. Fortunately, 0.01 seems to perform well in most cases.
@@ Appendix F Experimental setting/details
Appendix G Additional Experiments
To further validate the robustness and scientific utility of our proposed dataset and model, we conducted a series of additional experiments in response to reviewer feedback. These experiments evaluate the model’s generalization capabilities across different domains, its performance on downstream scientific tasks, its robustness to noise, and its computational efficiency.
G.1 Generalization and Robustness Analysis
Cross-Filter Generalization: To test the model’s performance on data from different instrumental filters, we evaluated our F814W-trained model on test sets from the F606w and F435w filters of the Hubble Space Telescope (HST). As shown in Tab. 10, while there is a performance drop as the filter domain shifts further from the training domain (F814W), the model maintains strong performance, demonstrating satisfactory generalization capabilities. The F606w filter, being spectrally closer to F814W, yields better results than the more distant F435w filter, confirming that domain similarity influences performance.
Table 10: Cross-filter generalization performance of the FISR model trained on the F814W filter.
Metric F435w F606w F814w (In-Domain)
PSNR 35.9192 36.3522 37.8779
SSIM 0.7305 0.7667 0.8311
Flux Error 0.9193 0.8242 0.5739
Robustness to Noise: We evaluated FISR’s robustness by introducing random Poisson noise to each image during inference, simulating realistic observational noise. The results in Tab. 11 show that FISR maintains its state-of-the-art performance, achieving the best results across all metrics compared to other methods under noisy conditions.
Table 11: Performance comparison under Poisson noise injection during inference. Best results are in bold.
Method Bicubic EDSR SwinIR RCAN HAT RealESRGAN FISR (Ours)
PSNR 28.9823 34.8191 36.3945 35.2918 36.8342 35.8098 36.7803
SSIM 0.6825 0.7684 0.7883 0.7848 0.7743 0.7852 0.7888
Flux Error 4.7889 1.5682 1.1433 1.1993 1.1943 6.9292 1.1025
Cross-Dataset Evaluation: Although direct evaluation is challenging due to differences in data units (STAR uses scientific counts, while AstroSR uses RGB), we re-trained our FISR model on the AstroSR dataset. Tab. 12 demonstrates that our method outperforms the original baseline models reported in the AstroSR paper, showcasing its architectural effectiveness on different data types.
Table 12: Performance comparison on the AstroSR dataset after re-training. Best results are in bold.
Method Bicubic EDSR RCAN ENLCA SRGAN FISR (Ours)
PSNR 17.7714 23.2168 23.6082 23.4267 23.0039 24.0211
SSIM 0.1686 0.3910 0.3966 0.3963 0.3854 0.4025
Flux Error 233.2564 50.5872 61.3863 59.1659 42.3078 33.2331
G.2 Evaluation on Downstream Scientific Tasks
To quantify the practical impact of our super-resolution model on real-world scientific analysis, we evaluated its performance on four representative downstream astronomical tasks. These experiments are designed to demonstrate that improvements in standard metrics like PSNR, SSIM, and our proposed Flux Error (FE) directly translate to higher fidelity in scientific measurements. The methodologies and results for these tasks are detailed below, with a final comparative summary in Table 13.
G.3 Evaluation on Downstream Scientific Tasks
To quantify the practical impact of our super-resolution model on real-world scientific analysis, we evaluated its performance on two representative downstream astronomical tasks. These experiments are designed to demonstrate that improvements in standard metrics and our proposed Flux Error (FE) directly translate to higher fidelity in scientific measurements. The methodologies and results for these tasks are detailed below.
Object Detection Sensitivity: The ability to detect faint objects is fundamental to astronomical surveys, determining the depth and completeness of celestial catalogs. An effective SR model should enhance faint sources, thereby improving detection sensitivity. In our experiment, we performed bipartite matching between sources detected in the predicted images and the ground-truth catalog, with a match considered successful if the spatial distance was within 2 pixels. The sensitivity was quantified using the Recall metric. Our FISR model achieves a high recall of 81.47%, indicating strong performance in identifying celestial objects.
Distance Estimation: Accurately measuring the distances to celestial objects is a cornerstone of cosmology, combining both object detection and precise photometry. To evaluate this, we used the successfully matched object pairs from the detection task. We converted each object’s flux to an apparent magnitude ($m$) and then applied the distance modulus formula, $d=10^{(m-M+5)/5}$, to estimate the distance ($d$) in megaparsecs (MPC), assuming a constant absolute magnitude ($M$) of 4.83 (typical for Sun-like stars). The accuracy was evaluated by the Mean Absolute Error (MAE) between the predicted and ground-truth distances, with the results shown in Table 13.
Table 13: Evaluation on the downstream task of distance estimation. Lower values indicate better performance. Best results are in bold.
Metric Bicubic SwinIR EDSR RCAN HAT R-ESRGAN FISR
Distance MAE (MPC) 6.82E+03 5.37E+03 6.44E+03 5.61E+03 4.44E+03 4.89E+03 4.12E+03
Appendix H More Ablation Studies on FGG Module
We performed ablation studies to analyze the sensitivity of the Flux Guidance Generation (FGG) module.
Kernel Choice in FGG: We tested alternative kernels (Airy, and a random mix of Gaussian/Airy) for rendering the flux map. Tab. 14 shows that performance remains stable across different kernel choices, suggesting that the module’s primary function is to provide a spatial prior for flux information, rather than depending on a specific kernel formulation.
Table 14: Ablation study on the kernel choice within the FGG module.
Kernel Type PSNR SSIM Flux Error
Gaussian 37.8779 0.8311 0.5739
Airy 37.6988 0.8305 0.5664
Gaussian/Airy (Random) 37.8186 0.8311 0.5726
Sensitivity to Detection Errors: To assess FGG’s robustness, we introduced noisy detections by lowering the source detection threshold, resulting in twice the number of sources, including many false positives. As seen in Tab. 15, while performance degrades slightly, FISR remains robust and achieves results comparable to the model trained with clean detections. This indicates that the model’s performance does not solely depend on the precision of the FGG’s input.
Table 15: Performance of FISR with clean versus noisy source detections in the FGG module.
Detection Quality PSNR SSIM Flux Error
Clean Detections 37.8779 0.8311 0.5739
Noisy Detections 37.3176 0.8275 0.6872
Appendix I Computational Efficiency
We measured the single-image inference time for all compared methods. The results in Tab. 16 show that FISR is computationally efficient, with an inference time comparable to other high-performing transformer-based models like SwinIR.
Table 16: Inference time per image (in seconds) for various SR methods.
Method Bicubic EDSR SwinIR RCAN HAT RealESRGAN FISR (Ours)
Time (s) 0.0014 0.1908 0.1088 0.1237 0.6747 0.0995 0.1698
Appendix J Limitations and future work
@@ Appendix I Computational Efficiency
Table 17: Ablation study on the penalty factor $\lambda$ ($\times 2$ on Gaussian PSF + Airy PSF).
@@
Appendix K More Visualizations of the STAR Dataset
To further illustrate the unique characteristics and scale of the STAR benchmark, this section provides additional visualizations of the source data. We present examples of the original, full-frame observational images from the Hubble Space Telescope (HST) WFC/ACS instrument, which constitute the raw data prior to the patch subdivision process for model training 7.
These wide-field views underscore a core advantage of STAR over previous object-centric datasets. Instead of focusing on isolated, cropped targets, our dataset provides a holistic view of extensive celestial regions, preserving the crucial spatial context and inter-object relationships (e.g., cross-object interaction, weak lensing). Furthermore, we showcase a gallery of selected image patches to highlight the rich diversity within STAR 8. These examples span a wide range of astronomical environments, from dense, crowded stellar fields and sparsely populated regions to complex nebulae and fields containing multiple galaxies. Collectively, these visualizations reinforce the value of STAR as a comprehensive and physically faithful benchmark for advancing astronomical super-resolution research.
Figure 7: Examples of the original wide-field raw data from the HST WFC/ACS survey, which form the basis of the STAR dataset.
Figure 8: A selection of patches from the STAR dataset, showcasing its diversity. The examples include crowded stellar fields, regions with interacting galaxies, and complex nebulae, demonstrating the variety of astronomical scenes available for training robust models.